Your piggy bank contains $19.75 in dimes and quarters. There are 100 coins in all. How many dimes are there?

Answer by CharlesG2(834)

Your piggy bank contains $19.75 in dimes and quarters. There are 100 coins in all. How many dimes are there?

D = number of dimes, Q = number of quarters
D + Q = 100
0.10D + 0.25Q = 19.75 or 10D + 25Q = 1975

D = 100 - Q
10(100 - Q) + 25Q = 1975
1000 - 10Q + 25Q = 1975
-10Q + 25Q = 975
15Q = 975
Q = 65 quarters
D = 100 - 65 = 35 dimes
check: 10(35) + 25(65) = 350 + 1625 = 1975